Let Z ∼ N(0, 1).

Find a constant c for which P(Zc) = 0.1587. Round the answer to two decimal places.

Find a constant c for which P(cZ ≤ 0) = 0.4772. Round the answer to two decimal places.

Find a constant c for which P(−cZc) = 0.8664. Round the answer to two decimal places.

Find a constant c for which P(0 ≤ Zc) = 0.2967. Round the answer to two decimal places.

Find a constant c for which P(|Z| ≥ c) = 0.1470. Round the answer to two decimal places.
Answer #1

-) Answer: To find c when z N(o, 1) p(2> c) = 0.1587 since p(271) = 0.1587 comparing (1) and (2), | c= 1.00 e) 2 >N(0,1) HA00Pi-CS ZSC) = 0.8664 >> 1-2p(Z>c) = 0.8664_ > 1-0, 8664 - 2P(2>C). => 0.1336 = 2 P(2@7c) 32 _p(2>c) 0.9336/2 0) P(Z>c) - 0.066vl P(1217,C) = 0.1470 > I-p(121<C) = 0.1470 l-0.1470 = P(12]<c) => P(Z <C) = 0.853 >> => - PC-c<Zcc) - 0.853 1-2pcZ>c) = 0.85